OpenStudy (anonymous):

Given the linear equation x + y = 5, find the slope of its graph. - 5 - 1 1 5

hero (hero):

To find the slope of any linear equation, write it in the form y = mx + b. In this case, you have x + y = 5. To find the slope re-write it in slope intercept form as follows: x + y = 5 y = 5 - x y = -x + 5 The last form y = -x + 5 is in the form y = mx + b, where m = -1. Therefore, the slope is -1.

OpenStudy (anonymous):

your my "hero".... haha jk, but thank you for explaining it :)